Derive the single particle partition function for an ideal monatomic gas consisting of N identical particles occupying a volume V. Obtain expression for the entropy and pressure of this system.

1
Expert's answer
2022-03-14T09:45:21-0400

Answer

The dispersion ratio of ultrarelativistic particles has the form and also partition function of a monatomic classical gas.


"C_p=C_v+1=4"

Due to the thermal equation of state of an ideal gas p = θ / v and cv = const, the equation of the adiabatic process has the form as beolw.

"PV^{\gamma}=constant."

Where

"\gamma=4/3"